# A Semantics-Aware Hierarchical Self-Supervised Approach to Classification of Remote Sensing Images

### v1

@@ V-A Case I: VHR Scene Classification
TABLE I: Comparison of the Overall Accuracy obtained by different hierarchical methodologies with different backbone architectures on the NWPU-RESISC45 dataset. The best results are reported in bold, second best are underlined. The related standard deviation is reported in brackets.
NWPU-RESISC45
Backbone Methodology OA (%)
LWGANet L2 [35] Baseline 92.50 (0.31)
HAF [18] 93.05 (0.29)
HRN [20] 93.01 (0.11)
SAHC (proposed) 94.69 (0.18)
MobileNet V2 2.5× [34] Baseline 93.70 (0.30)
HAF 93.67 (0.28)
HRN 93.58 (0.16)
SAHC (proposed) 94.19 (0.33)
EfficientformerV2 S2 [33] Baseline 90.41 (0.38)
HAF 90.76 (0.46)
HRN 91.16 (0.44)
SAHC (proposed) 91.72 (0.14)
Table I shows the OAs of the LWGANet, MobileNet, and EfficientFormer on the NWPU-RESISC45 dataset. The proposed SAHC achieves the best results, showing a higher OA% with all the different backbones. The HAF and HRN hierarchical modules show improvements only in the case of the transformer-based backbone (i.e., EfficientFormer) and the hybrid backbone (i.e., LWGANet), while the baseline approach without hierarchical modules performs better with the CNN-based approach (i.e., MobileNet). In contrast, the proposed SAHC shows an increment in OA% of $\sim 1.2\%$ compared to the baseline approach with both the transformer-based and the CNN-based architectures, while exhibits a greater improvement when considering the hybrid backbone, with an additional $\sim 2.2\%$ increase in mF1%. These findings highlight the contribution of the hierarchical module. Indeed, its application to the hybrid architecture results in higher performance compared to its application to the CNN model. This result is particularly interesting given that the baseline CNN model outperforms the LWGANet baseline. It is also worth noting that, in most cases, the SAHC shows a smaller variability compared to its competitors, ensuring stabler results across different random initializations of the architecture weights.
@@ V-B Case II: MS SITS Semantic Segmentation
TABLE II: Comparisons of the accuracies obtained by different hierarchical methodologies with different backbone architectures on the ELU and HRLC-CCI datasets. The best results are reported in bold, second best are underlined. The related standard deviation is reported in brackets.
ELU Dataset HRLC-CCI Dataset
Backbone
Methodology
mF1 (%)
OA (%)
mF1 (%)
OA (%)
Swin Transformer [37] Baseline (no hierarchy) 33.42 (0.44) 68.10 (0.65) 64.03 (1.06) 83.30 (1.22)
HAF [18] 32.97 (0.57) 68.15 (0.56) 64.26 (0.41) 83.35 (0.68)
HRN [20] 30.48 (0.86) 67.72 (1.98) 63.80 (0.38) 83.05 (0.49)
SAHC (proposed) 34.57 (0.12) 71.25 (0.25) 66.49 (0.34) 84.75 (0.47)
Transformer [36] Baseline (no hierarchy) 18.77 (0.29) 54.79 (0.87) 61.14 (0.62) 81.74 (0.72)
HAF 19.04 (0.28) 55.69 (0.88) 61.12 (0.81) 81.37 (1.03)
HRN 20.97 (0.58) 57.48 (2.31) 60.46 (1.05) 81.80 (0.60)
SAHC (proposed) 25.63 (0.41) 65.94 (0.74) 64.07 (0.45) 83.30 (0.44)
ConvLSTM [38] Baseline (no hierarchy) 25.82 (0.95) 62.59 (0.70) 65.19 (0.77) 84.32 (0.33)
HAF 26.11 (0.92) 62.40 (0.57) 65.64 (0.84) 84.42 (0.68)
HRN 25.53 (1.03) 62.36 (1.47) 65.18 (0.31) 84.19 (0.58)
SAHC (proposed) 26.48 (0.35) 66.74 (0.42) 65.38 (0.40) 85.08 (0.28)
The second study case analyzes the semantic segmentation task, and is characterized by two different datasets with varying hierarchical complexity and differing label quality. The ELU dataset employs weak labels due to its focus on land use rather than land cover. Conversely, the HRLC-CCI dataset is designed for pixel-based LC classification, providing more precise and location-specific labels. Table II shows the performance of the three considered backbones on both the ELU and the HRLC-CCI dataset in terms of mF1% and OA%. The proposed approach outperforms the other methods on the ELU dataset, which exhibits stronger hierarchy and weaker labels than the HRLC-CCI dataset. Notably, the pixel-based transformer without contextual information shows an improvement of about $7\%$ in terms of mF1% compared to the non-hierarchical baseline, while it also outperforms the other hierarchical approaches. In the case of the hybrid Swin Transformer and the ConvLSTM, the performance gap tends to decrease, while the adoption of the hierarchy and the consensus approach still shows an improvement. This result underlines the robustness of the proposed method to noisy labels, as the trained hierarchical matrices can learn soft hierarchical relationships directly from the data. In contrast, the compared methodologies rely on a hard hierarchy definitions, which might lead to suboptimal results in the case of labels that do not accurately represent the observed samples. The same behavior can be observed when analyzing the results on the HRLC-CCI dataset, where the hierarchy becomes weaker, but the real targets are better represented by the labels at the pixel level. In this case, the proposed approach performs similarly to the other methodologies, as the hierarchy is less noisy and the mapping more robust. Nonetheless, SAHC still outperforms the other methodologies, showing good generalization across the different backbones and scenarios. In our experiments, SAHC performed worse than HAF in terms of mF1% score only in the case of the ConvLSTM backbone, although the performance remains comparable. However, the proposed approach still shows a higher OA%, with reduced scores variability across different seeds. Lastly, Fig. 4 shows a visual comparison of the different hierarchical methodologies considering the HRLC-CCI dataset and the Swin Transformer backbone. From the qualitative results, one can see that the SAHC approach better represents the “Tree” classes, also showing a lower pixel-level noise.
Figure 4: Example of qualitative results at the fine-grained level on the HRLC-CCI dataset obtained by the considered methods using the Swin Transformer backbone.
V-C Ablation Study
TABLE III: Ablation study on the Swin backbone with the ELU dataset. The classification metrics are reported with the related standard deviation in brackets. The best results are reported in bold, second best are underlined.
Multi-level
Classifiers Hierarchical
Consensus
(Training) Hierarchical
Consensus
(Inference) mF1 (%)
OA (%)
33.42 (0.44) 68.10 (0.65)
33.48 (0.59) 68.10 (1.06)
34.10 (0.41) 67.83 (1.00)
33.48 (0.37) 68.39 (0.33)
34.57 (0.12) 71.25 (0.25)
@@ V-C Ablation Study
The performance of each component in the proposed hierarchical approach is further analyzed in the ablation study presented in Table V-C. In this study, we evaluate the Swin backbone across five distinct configurations on the ELU dataset:
@@
    Baseline: the baseline approach with no multi-level classifiers;
@@
    Multi-level: the baseline approach augmented with additional classification heads for the coarse-grained and intermediate classes;
@@
    Multi-level with HC inference: the multi-level approach with consensus applied solely during inference (considering user-defined hierarchies);
@@
    Multi-level with $\mathcal{L}_{HC}$ training: the multi-level approach trained with the self-supervised consensus loss, excluding consensus-based classification during inference;
@@
    SAHC: the proposed approach, which incorporates both the self-supervised consensus loss and consensus inference at multiple levels.
@@
Figure 5: Heatmap of the estimated hierarchy log-joint matrices considering the Swin Transformer backbone on the ELU dataset from fine-grained classes to intermediate classes. The blue and green rectangles identify the expected hierarchical aggregation of the fine-grained classes at the intermediate and coarse levels, respectively.
The table results indicate that incorporating multi-level classifiers into the baseline does not improve the overall performance. In fact, it leads to increased variability in the results, showing a higher variance and making the outcomes more unstable. However, the introduction of the HC inference improves the representation of fine-grained classes in terms of mF1%, also when considering the user-defined hierarchy in the multi-level approach with HC inference. On the contrary, the multi-level approach with $\mathcal{L}_{HC}$ training does not improve the performance, as the JSD acts on the consensus across the layers and not on the fine-grained classification itself. This is instead addressed by the proposed SAHC, which combines the self-supervised consensus loss with consensus inference at multiple hierarchy levels. By allowing the trained matrices to learn the relationships from the data, the hierarchical consensus exhibits a sharp improvement in fine-grained classification, outperforming all other configurations. It is also important to note that the introduction of the JSD and the hierarchical consensus in inference leads to lower variance values, showing a more consistent fine-grained classification across different weights initializations compared to the other approaches.
V-D Hierarchy Projection Matrices Analysis
Figure 6: Heatmap of the estimated hierarchy log-joint matrices considering the Swin Transformer backbone on the HRLC-CCI dataset from fine-grained classes to intermediate classes. The blue and green rectangles identify the expected hierarchical aggregation of the fine-grained classes at the intermediate and coarse levels, respectively.
To better visualize the impact of the hierarchy projection matrices on classification results, we analyze the estimated hierarchy projection matrices of the Swin Transformer on the ELU and HRLC-CCI datasets. The visualization on the NWPU dataset is not reported as the hierarchical mappings show a trivial result, with perfect mapping of the defined hierarchical levels (mostly due to the simpler classification task).
Figure 5 presents a heatmap of the log-joint matrix between intermediate and fine-grained levels, highlighting the hierarchical structure captured by the proposed methodology on the ELU dataset. Note that the hierarchy projection matrix was uniformly initialized before training, with no prior knowledge of the user-defined hierarchical structure of the problem. The results show that the proposed approach effectively learns the hierarchy in the dataset. Specifically, the blue rectangles in the heatmap correspond to the user-defined expected mappings, from fine-grained classes to intermediate classes. Here, the weights associated to these mappings are highlighted, confirming that the data-driven estimated mappings are correctly following the user-defined hierarchies. Then, green rectangles indicate distant mappings to coarse-grained classes, revealing a similar pattern even though the matrix focuses on mappings between fine-grained and intermediate levels. It is important to note that separate hierarchy joint matrices are trained for the different hierarchical mappings in the proposed approach.
Similarly, Figure 6 depicts the heatmap associated with the HRLC-CCI dataset. As observed in the ELU case, the proposed methodology correctly extracts the hierarchy within the problem analyzed, emphasized by the blue rectangle showing the intermediate aggregation. However, the mappings appear more dispersed compared to the more complex ELU scenario, mostly due to the simpler hierarchical constraint leading to a smaller impact. Nonetheless, the trainable log-joint matrices allow soft-hierarchical mappings that were not considered in the initial user-defined hierarchy. One example can be observed in the weights associated to “Grassland vegetation aquatic or regularly flooded areas”, which shows an active semantic relationship with the “Open water” class, which is not mapped by the imposed user-defined hierarchy. This mapping highlights the importance of a data-driven approach, in contrast with the superimposition of a rigid hierarchical structures.

### v2

@@ V-A Case I: VHR Scene Classification
The performance of the proposed SAHC method on the NWPU-RESISC45 dataset is summarized in Table I, providing a comparative analysis against state-of-the-art methodologies and the non-hierarchical baseline. Given the balanced class distribution characteristic of the NWPU-RESISC45 dataset, the OA and mF1 are aligned across all evaluated models. From the table, one can see that the integration of hierarchical information consistently improves classification performance over the baseline. Indeed, the majority of the evaluated hierarchical architectures outperform their non-hierarchical counterpart (with the exception of Hierarchical Contrastive Learning (HCL)), confirming that the exploitation of semantic relationships allows for a more robust distinction between fine-grained RS classes. Among the compared methodologies, the proposed SAHC achieves the highest performances at the fine-grained level, with an OA of 96.6396.63% and an mF1 of 96.6396.63%. Notably, the good performance of SAHC extends across the entire taxonomic hierarchy. At the intermediate and coarse-grained levels, the proposed method achieves OAs of 97.5697.56% and 97.9897.98%, respectively, consistently outperforming the other frameworks. The second-best overall performance is obtained by the HRN method, which reaches a fine-grained OA of 96.4896.48%. However, this comparison also highlights the computational overhead introduced by the different architectures. The HRN method achieves its high accuracy at the cost of architectural efficiency. Indeed, the introduction of the Granularity-Specific Block (GSB) leads to a significant increase in parameters and computational overload. In contrast, the proposed SAHC maintains the same computational footprint as the baseline architecture. This efficiency is shared by other methods such as Adaptive Label Granularity Selection (ALGS), HAF, and HiCervix, where the bottleneck remains the backbone itself. In the specific case of SAHC, the architectural overhead is due to the three auxiliary heads and their associated projection matrices, both of which are negligible relative to the dimension of the backbone. When focusing strictly on architectures that do not introduce additional complexity, the ALGS method provides the highest performance after SAHC, though it remains lower in terms of OA.
@@ V-B Case II: MS SITS Semantic Segmentation
TABLE II: Comparison of the Accuracies obtained by different hierarchical methodologies on the ELU dataset. The best results are reported in bold, second best are underlined
Metric Hierarchy Baseline [22] HRN [8] HAF [15] ALGS [7] HCL [9] MGM [31] HiCervix [4] SAHC
OA (%) Fine-grained 36.39 36.41 36.63 35.90 37.26 35.42 36.92 38.08
Intermediate 46.60 41.58 46.21 48.81 45.99 47.91 49.68
Coarse-grained 57.13 57.89 56.43 59.58 56.30 58.27 59.90
Macro 76.34 78.82 75.49 78.36 75.44 77.39 78.35
Top-3 (%) Fine-grained 56.48 57.57 55.86 54.75 56.94 55.83 56.98 58.96
Intermediate 71.01 63.61 69.27 71.59 70.13 71.31 73.48
Coarse-grained 82.44 81.15 80.90 82.76 81.66 82.29 84.31
Macro 96.77 96.36 97.10 97.30 96.99 97.50 97.23
Top-5 (%) Fine-grained 64.57 66.61 64.21 63.11 65.54 64.89 65.48 67.84
Intermediate 80.68 74.55 78.72 80.60 79.92 80.66 82.80
Coarse-grained 90.67 88.88 90.07 90.69 90.18 90.93 91.98
Macro -— -—
mF1 (%) Fine-grained 32.02 32.38 32.12 31.63 33.27 31.60 32.59 34.25
Intermediate 39.25 27.92 38.60 41.03 39.19 39.86 42.59
Coarse-grained 52.64 46.42 51.44 54.80 51.98 53.23 55.34
Macro 66.96 70.64 67.12 69.47 66.48 68.95 69.12
Params (M) 4.7 16.2 4.8 4.8 5.8 16.2 4.7 4.8
FLOP (T) 2.22 2.38 2.22 2.22 2.41 2.32 2.22 2.22
To further evaluate the robustness of the proposed framework, an extensive analysis was conducted on the highly complex ELU dataset. Unlike the previous scenario, this dataset consists of lower-resolution multispectral images (10m10m/pixel) mapped to a complex taxonomy of $90$ classes distributed across four hierarchical levels. The results of this analysis are reported in Table II. The complexity of the ELU dataset, combined with the potential for semantic noise or inconsistencies in the CLC LU, makes this scenario particularly challenging. Standard hierarchical models that incorporate a rigid taxonomy might struggle to adapt to these inconsistencies. The proposed SAHC mitigates this limitation by leveraging a data-driven taxonomy, allowing the architecture to dynamically adapt even when the predefined hierarchy is not perfectly aligned. This advantage can be observed in the OA and mF1 shown in the Table. Indeed, SAHC achieves the best performance at the fine-grained level with an OA of 38.0838.08% and an mF1 of 34.2534.25%. These results are also confirmed at the intermediate (OA 49.6849.68%, mF1 42.5942.59%) and coarse-grained levels (OA 59.9059.90%, mF1 55.3455.34%), consistently better performing with respect to other state-of-the-art methodologies like HiCervix and HCL. At the macro level, SAHC demonstrates competitive performance, ranking after only to the HCL and HAF methods by a marginal gap.
To further evaluate the ability of the models to capture the semantic relationships, we analyzed the Top-3 and Top-5 accuracy metrics across all the levels of the hierarchy (see Table II). These metrics indicate whether the ground-truth class is among the highest-confidence predictions of the model, measuring the capacity of the network to maintain semantic proximity even when the absolute Top-1 prediction is incorrect. The proposed SAHC approach demonstrates the highest performance across almost all granularity levels. Specifically, SAHC achieves the highest Top-3 accuracies for the fine-grained, intermediate, and coarse-grained levels, reaching a 58.9658.96%, 73.4873.48%, and 84.3184.31%, respectively. This represents a large improvement against the HRN, which ranked second in the VHR case, showing a margin of 1.391.39% at fine-grained level, while also outperforming the RS-oriented competitor HCL by 1.891.89% and 1.551.55% in the intermediate and coarse levels, respectively. At the macro level, Top-3 performance saturates for all evaluated models. In this aggregated space, SAHC remains robust (97.2397.23%), performing comparably to HCL (97.3097.30%) and HiCervix (97.5097.50%). Because the total number of target classes at the macro level is small, Top-3 classification becomes a less discriminative metric. For the Top-5 accuracy, SAHC maintains the overall best performance, outperforming the baseline and the compared methodologies across the evaluated granularities. At the fine-grained level, SAHC achieves 67.8467.84%, surpassing the HRN by 1.231.23%. This performance gap widens at the intermediate level, where SAHC reaches 82.8082.80% compared to HRN 80.6880.68%. At the coarse-grained level, SAHC achieves 91.9891.98%, outperforming the HiCervix by 1.051.05%. Note that Top-5 metrics are omitted at the macro level, as the limited number of classes at this highest taxonomy level renders the Top-5 metric uninformative. The improvement observed in the Top-3 and Top-5 accuracies can be attributed to the design of the SAHC architecture. When the network struggles to isolate the exact fine-grained class, the hierarchical consensus mechanism, which ensures probabilistic alignment across different hierarchical levels, guarantees that the subsequent highest-confidence predictions belong to semantically related classes within the same hierarchical branch. This minimizes the divergence between the projected predictions and the overall committee consensus, reducing severe misclassifications.
Figure 5: Qualitative example of the predictions at the fine-grained level on the ELU dataset obtained by the considered hierarchical methodologies using the ResNet50 backbone: (a) true color (RGB) monthly composite for August 2020, (b) CLC reference map, (c) non-hierarchical baseline (d) HRN method, (e) HAF method, (f) ALGS method, (g) HCL method, (h) MGM method, (i) HiCervix method, and (j) the proposed SAHC method.
Finally, consistent with the previous analysis, SAHC achieves these state-of-the-art results without sacrificing computational efficiency. While architectures like HRN and multi-granularity module (MGM) impact on the computational efficiency, SAHC keeps the computational footprint nearly identical to the baseline model.
Lastly, Fig. 5 provides a visual comparison of the evaluated hierarchical methodologies on a small portion of the ELU dataset. From the qualitative results, one can see that the SAHC approach retrieves the most spatially-consistent land-cover maps, characterized by a significant reduction in classification artifacts compared to the competing approaches. This is particularly evident in the southern sector of the study area. Indeed, while in this specific case most networks misclassify “transitional woodland/shrub” as “inactive quarries” due to sparse vegetation signatures, SAHC correctly identifies the vegetative cover, showing robustness in classification. These advantages are further evident in Fig. 6. From a broader perspective, scaling a RS classification problem to an extensive taxonomy of $90$ classes inherently introduces severe spectral confusion and extreme class imbalance. In this context, the observed improvements in the mF1 score are significant. Indeed, they demonstrate that the proposed SAHC mitigates the ambiguities typical of such label space, maintaining robust and granular predictions where traditional approaches fail.
Figure 6: Qualitative example of the predictions at the fine-grained level on the ELU dataset obtained by the considered hierarchical methodologies using the ResNet50 backbone: (a) true color (RGB) monthly composite for August 2020, (b) CLC reference map, (c) non-hierarchical baseline (d) HRN method, (e) HAF method, (f) ALGS method, (g) HCL method, (h) MGM method, (i) HiCervix method, and (j) the proposed SAHC method.
V-C Sensitivity Analysis
Figure 7: Sensitivity analysis of the hierarchical weights $\lambda^{h}$ on the VHR dataset. The performance surface is visualized as a probability simplex, where each point represents a weight combination for the three hierarchical levels sampled at a step size of $0.1$. The simplex heatmap shows the optimal configuration as $\lambda^{1}=0.5$, $\lambda^{2}=0.2$, $\lambda^{3}=0.3$.
To evaluate the robustness of the proposed hierarchical architecture, we conducted a sensitivity analysis on the VHR dataset to assess how model performance varies with the selection of different granularity weights $\lambda^{h}$. The search space was constrained such that the weights sum to one ($\sum_{h=1}^{H}\lambda^{h}=1$), with each $\lambda^{h}$ sampled using a step size of $0.1$. Since the architecture adopts three hierarchical levels, the resulting parameter space can be visualized as a probability simplex that illustrates the various weight combinations across the different granularities. Fig. 7 illustrates a heatmap reporting the performance across all evaluated configurations of the three layer weights. As demonstrated in the figure, the optimal combination is achieved at $\lambda^{1}=0.5$, $\lambda^{2}=0.2$, $\lambda^{3}=0.3$. This optimal configuration justifies the hypothesis that priority should be allocated to both the fine-grained classification task and the coarsest class. Nevertheless, the analysis also reveals the presence of several local maxima across the performance surface, which are generally skewed toward configurations in favor of the finest-level weight.
V-D Ablation Study
TABLE III: Ablation study on the ResNet50 backbone with the
NWPU-RESISC45 dataset. The best results are
in bold, second best are underlined
Multi-Heads $L_{HC}$ $\hat{\mathbf{p}}_{HC}$ $L^{HC}_{CCE}$ Hierarchy OA (%) mF1 (%)
Fine 96.08 96.08
Inter.
Coarse
Fine 96.44 96.43
Inter. 97.41 96.68
Coarse 97.92 97.58
Fine 96.49 96.48
Inter. 97.44 96.68
Coarse 97.94 97.57
Fine 96.52 96.53
Inter. 97.41 96.64
Coarse 97.76 97.33
Fine 96.49 96.51
Inter. 97.48 96.71
Coarse 97.90 97.48
Fine 96.63 96.63
Inter. 97.54 96.86
Coarse 97.97 97.66
@@ V-D Ablation Study
The performance contribution of each component in the proposed hierarchical approach is analyzed in the ablation study presented in Table III. In this study, we evaluate the ResNet50 backbone across six different hierarchical configurations on the NWPU-RESISC45 dataset:
@@
    Baseline: the standard approach with no multi-level classifiers;
@@
    Multi-level: the baseline approach modified with additional classification heads for the coarse-grained and intermediate classes.
@@
    Multi-level with HC inference $\hat{\mathbf{p}}_{HC}$: the multi-level approach with consensus applied solely during inference (considering user-defined hierarchies).
@@
    Multi-level with $\mathcal{L}_{HC}$ training: the multi-level approach trained with the self-consistent consensus loss, excluding consensus-based classification during inference.
@@
    Multi-level with $\mathcal{L}_{HC}$ training and HC inference $\hat{\mathbf{p}}_{HC}$: the multi-level approach combining both the self-supervised consensus training and consensus-based classification during inference.
  6. 6.
@@
    SAHC: the proposed full approach, which incorporates the self-supervised consensus loss, consensus inference, and the supervised consensus loss ($\mathcal{L}^{HC}_{CCE}$) at multiple levels.
As observed in Table III, the integration of the proposed modules consistently enhances the performance of the network. The transition from the baseline to the simple multi-level architecture shows an increase of the fine-grained OA from 96.0896.08% to 96.4496.44%. This demonstrates that simply forcing the network to learn coarse and intermediate weighted representations provides regularization for the fine-grained task. Introducing the HC inference constraint or the self-consistent loss individually provides further performance gains. Notably, using only the self-consistent consensus training leads to good fine-grained OA of 96.5296.52% and an mF1 of 96.5396.53%, suggesting that enforcing hierarchical consistency during the learning phase is effective. Interestingly, combining both modules creates a trade-off, reducing marginally the fine-grained accuracy while improving the intermediate OA to 97.4897.48%. Finally, the introduction of the supervised consensus loss in the complete SAHC framework provides the best trade-off between training and inference constraints. The full SAHC configuration achieves the highest metrics across all hierarchical levels, achieving the best fine-grained OA of 96.6396.63%. This confirms that the cooperation between self-consistent alignment, structured inference, and supervised hierarchical correction maximises the discrimination capabilities of the network.
V-E Hierarchy Projection Matrices Analysis
@@
Figure 8: Heatmap of the estimated joint probability matrix $\mathbf{J}^{h_{2},h_{1}}$ considering the ResNet50 backbone on the ELU dataset from fine-grained classes to intermediate classes. The cyan rectangles identify the expected hierarchical aggregation of the fine-grained classes at the intermediate level.
To better visualize the impact of the hierarchical projection matrices on classification results, we analyze the estimated joint probability matrix $\mathbf{J}^{h_{2},h_{1}}$ representing the learned hierarchy between the fine-grained and intermediate levels for the ELU dataset. The corresponding visualization on the NWPU dataset is omitted, as its simpler classification task leads to an expected result where the learned mappings perfectly align with the predefined hierarchical levels. Figure 8 presents a heatmap of this joint probability matrix, illustrating the hierarchical structure captured by the proposed SAHC methodology on the ELU dataset. The visualization confirms that the proposed approach effectively learns the hierarchy in the dataset, correctly identifying the semantic aggregations between the levels. Specifically, the cyan rectangles in the heatmap correspond to the user-defined expected mappings, from fine-grained classes to intermediate classes. The higher weights within these regions confirm that the data-driven estimated mappings successfully respect the intended user-defined taxonomy. In addition to these rigid hierarchies, the network also captures important semantic relationships between different coarser groups. Indeed, the blue rectangles indicate cross-groups relationships related to to coarser-grained classes. Because separate joint matrices are trained for different hierarchical transitions in the proposed architecture, the model is able to discern soft-hierarchical mappings that were strictly excluded from the initial user-defined hierarchy. For instance, the learned weights reveal an active semantic relationship between “water basins” and “aquaculture in the marine environment”, as well as between “continuous urban fabric” and “road networks”. These mappings highlight the importance of a data-driven approach. Indeed, rather than superimposing a rigid and potentially flawed hierarchical structure, the model dynamically adapts to the true semantic distributions present in the RS image.